the sum two numbers is 37. the sum of the samller number and 2 times the larger number is 61. find the numbers
Answer by checkley79(3341)

X+Y=37
X+2Y=61 SUBTRACT THE FIRST EQUATION.
-X-2=-37
----------
Y=24 ANS. FOR THE LARGER NUMBER.
X+24=37
X=37-24
X=13 ANS. FOR THE SMALLER NUMBER.
PROOF:
13+2*24=61
13+48=61
61=61